OpenStudy (anonymous):

a police ar traveling at 96.3 km/h is traveling west, chasig a motorist traveling at 72.6 km/h. if they are originally 250 m apart in what time interval will the police car overtake the motorist?

OpenStudy (anonymous):

Solve for t: 72.6*t + .25 = 96.3*t t (96.3 - 72.6) = .25 t = .25/(96.3 - 72.6) t =. 01055 hours This is about 37.98 seconds.

OpenStudy (anonymous):

wat kinematic equation/motion equation did u use?